# A Particle Algorithm for Mean-Field Variational Inference

### v2

@@ Abstract
Variational inference is a fast and scalable alternative to Markov chain Monte Carlo and has been widely applied to posterior inference tasks in statistics and machine learning. A traditional approach for implementing mean-field variational inference (MFVI) is coordinate ascent variational inference (CAVI), which relies crucially on parametric assumptions on complete conditionals. In this paper, we introduce a novel particle-based algorithm for mean-field variational inference, which we term PArticle VI (PAVI). Notably, our algorithm does not rely on parametric assumptions on complete conditionals, and it applies to the nonparametric setting. We provide non-asymptotic finite-particle convergence guarantee for our algorithm. To our knowledge, this is the first end-to-end guarantee for particle-based MFVI.
@@ 1 Introduction
Variational inference (VI) is a popular Bayesian inference method for approximating intractable posterior distributions (Jordan et al., 1999; Bishop, 2006; Wainwright et al., 2008; Blei et al., 2017). The premise of VI is to perform such approximation via a constrained optimization problem of the form
@@
where $\pi\in\mathcal{P}(\mathbb{R}^{n})$ is the posterior distribution, $\mathcal{Q}\subseteq\mathcal{P}(\mathbb{R}^{n})$ is some set of probability distributions, and $D_{\mathrm{KL}}(\cdot\|\cdot)$ is the Kullback–Leibler divergence. A common choice for $\mathcal{Q}$ is the set of product measures, i.e., $\mathcal{Q}=\big{\{}\prod_{i=1}^{n}q_{i},\;q_{i}\in\mathcal{P}(\mathbb{R})\big{\}}$, which is called the mean-field family. The corresponding variational problem for this choice of $\mathcal{Q}$ is called mean-field variational inference (MFVI). Compared to Markov chain Monte Carlo (MCMC) methods, which can have long burn-in periods, the optimization approach of VI allows for faster and more scalable inference in practice (Blei et al., 2017).
@@
Optimization problems over spaces of probability distributions can in theory be solved with a gradient flow by endowing the space of probability distributions with the 2-Wasserstein metric $\mathcal{W}_{2}$ and using a discrete stepping scheme (Ambrosio et al., 2005; Ambrosio and Savaré, 2007; Santambrogio, 2017). The use of $\mathcal{W}_{2}$–gradient flows to solve the variational problem (1.1) has led to developments in theoretical understanding of VI (Lambert et al., 2022; Yao and Yang, 2022; Diao et al., 2023; Lacker, 2023) as well as the unconstrained version of (1.1) whose solution is the exact posterior (Trillos and Sanz-Alonso, 2020). Furthermore, a seminal result (Jordan et al., 1998) introduces a close connection between $\mathcal{W}_{2}$–gradient flows and PDEs: the discrete iterative scheme that describes the $\mathcal{W}_{2}$–gradient flow converges, in the small step-size limit, to the solution of a Fokker–Planck equation (Jordan et al., 1998, Theorem 5.1). The solution to the Fokker–Planck equation, in turn, describes the law of a particle evolving according to a stochastic differential equation (SDE). This PDE-and-SDE characterization of the $\mathcal{W}_{2}$–gradient flow opens up avenues for computation: it is generally difficult to compute Wasserstein distances, and thus gradient flows in the metric space, but the corresponding PDE and SDE are easier to approximate. Furthermore, the PDE and SDE are more amenable to theoretical analysis and convergence guarantees.
@@
Since VI operates in a subset $\mathcal{Q}$ of the $\mathcal{W}_{2}$ space, efforts have been made to describe variants of VI—namely, VI with different choices of $\mathcal{Q}$—in terms of gradient flows. The difficulty lies in showing that a gradient flow constrained to a subspace of the $\mathcal{W}_{2}$ space is still a gradient flow. In other words, the challenge is to prove that the gradient flow is closed with respect to the subspace $\mathcal{Q}$. When this is shown, theoretical guarantees follow (Lambert et al., 2022; Yao and Yang, 2022; Lacker, 2023). In the case of MFVI, a recent paper (Lacker, 2023) shows that MFVI can be described by a $\mathcal{W}_{2}$–gradient flow constrained to the submanifold of product measures endowed with the $\mathcal{W}_{2}$ distance. The paper also shows that this gradient flow corresponds to a McKean–Vlasov equation obtained by an independent projection of the corresponding Langevin dynamics. The law of this McKean–Vlasov equation evolves according to a nonlinear Fokker–Planck equation, which is a system of coupled PDEs. This PDE-and-SDE description allows Lacker (2023) to specify convergence rates of the gradient flow to the solution of (1.1) as time goes to $\infty$.
@@
In this paper, we introduce a particle-based algorithm for MFVI, which we term PArticle VI (PAVI). The algorithm is built upon the SDE description of MFVI and stochastic approximation. Under standard regularity assumptions, we provide an upper bound on the 2-Wasserstein distance between the empirical measure of the particles at each iteration of Algorithm 2 and the solution of (1.1). Notably, our convergence guarantee is non-asymptotic and holds for any number of particles and total time duration. To our knowledge, this is the first end-to-end guarantee for particle-based MFVI.
The theoretical challenge lies not only in proving the convergence guarantees but also in designing appropriate time-discretization and particle approximation schemes for the McKean–Vlasov SDE. Other particle-based VI algorithms have been proposed in Yao and Yang (2022), but their theoretical results include error terms that do not diminish in the long-time limit and do not analyze the error stemming from the particle approximation. Proposing an implementable and efficient algorithm for the McKean–Vlasov SDE is not straightforward: as discussed in (Lacker, 2023, Section 2.3), the most obvious discretization does not contain a dissipative drift (which is needed for establishing uniform-in-time convergence rates).
Our algorithm not only has strong convergence guarantees (as discussed earlier) but is also nonparametric and flexible. The traditional algorithm for implementing MFVI, called coordinate ascent variational inference (CAVI), requires conditional conjugacy of the model parameters’ distributions (Bishop, 2006; Blei et al., 2017); in particular, the model’s complete conditionals must belong to exponential families. Our proposed method is both nonparametric and theoretically sound, like MCMC, while also being scalable and efficient thanks to the mean-field family of product measures, similarly to CAVI. On the mathematical side, we contribute novel results and techniques to the convergence analysis of McKean–Vlasov systems.
@@
We outline recent theoretical advances in VI based on gradient flows in the $\mathcal{W}_{2}$ space as follows. When $\mathcal{Q}$ is the set of Gaussians, the resulting form of VI is called Gaussian VI. The set of Gaussians endowed with the $\mathcal{W}_{2}$ distance is also called the Bures–Wasserstein (BW) space (Bures, 1969), which is a subset of the $\mathcal{W}_{2}$ space. By studying $\mathcal{W}_{2}$–gradient flows constrained to the BW space, Lambert et al. (2022) propose a stochastic gradient descent algorithm for Gaussian VI—which discretizes the gradient flow—and provide theoretical guarantees, and Diao et al. (2023) develop a forward-backward algorithm along with convergence rates for Gaussian VI. These works suggest extensions to the case when $\mathcal{Q}$ is the set of mixtures of Gaussians, which corresponds to an essentially nonparametric $\mathcal{Q}$. See Yi and Liu (2023); Huix et al. (2024) for further developments.
@@
In the mean-field case where $\mathcal{Q}$ consists of product distributions, Yao and Yang (2022) establish convergence guarantees for a coordinate-wise $\mathcal{W}_{2}$–gradient flow for MFVI in Bayesian models, and they propose two algorithms for practical implementation: a particle algorithm based on the Langevin SDE, and a function approximation method using neural networks. Their particle algorithm suffers from cumulative numerical errors that persist in the long-time limit (see (Yao and Yang, 2022, Lemma H.1 and Theorem D.2)), and they do not analyze the error stemming from the particle approximation. Lacker (2023) shows that MFVI can be described by a $\mathcal{W}_{2}$–gradient flow constrained to the submanifold of product distributions and establishes continuous-time convergence guarantees. Jiang et al. (2023) propose approximating the MFVI solution through optimization over polyhedral subsets of the Wasserstein space, and they introduce projected gradient descent algorithms with theoretical guarantees. Tran et al. (2023) also investigate particle MFVI; however, their theoretical analysis contains errors (see the footnote on (Jiang et al., 2023, Page 14)). Besides the above-mentioned algorithmic and computational developments for MFVI, there has also been recent work on the statistical accuracy of the solution to the MFVI problem for approximating the true posterior $\pi$. We refer the reader to, for example, Zhang and Gao (2020); Mukherjee and Sen (2022); Lacker et al. (2024) and the references therein for details.
@@
Our convergence criteria resemble those from the literature on Langevin Monte Carlo (LMC), see Dalalyan (2017b); Durmus and Moulines (2017); Cheng and Bartlett (2018); Durmus et al. (2019); Durmus and Moulines (2019); Balasubramanian et al. (2022). LMC simulates an SDE whose stationary distribution is proportional to $e^{-V(\cdot)}$, where $V$ is a given potential. Non-asymptotic theoretical guarantees for LMC are given under the condition that the potential $V$ is continuously differentiable, strongly convex, and has Lipschitz gradient; see Durmus and Moulines (2019) for convergence guarantees in the $\mathcal{W}_{2}$ distance. We refer the reader to Chewi (2023) for a survey on LMC.
@@
Our algorithm is built on the framework of gradient flows and particle algorithms. Gradient flows and particle algorithms have recently been applied to a wide range of problems in statistics and optimization. For example, Saeedi et al. (2017) propose a particle algorithm for approximating discrete distributions with the approximating family given by distributions with at most $K$ atoms (where $K$ is fixed). Cheng et al. (2024) propose particle algorithms for distribution approximation using generalized Wasserstein gradient flow of the Kullback-Leibler divergence. Their algorithms do not impose explicit specification of an approximating family $\mathcal{Q}$ and rely on neural networks to estimate certain vector fields used for updating. Their theoretical analysis relies on assumptions on neural network approximation accuracy and does not consider the error arising from particle approximation. We also refer the reader to Liu and Wang (2016); Gallouët and Monsaingeon (2017); Chizat and Bach (2018); Liu et al. (2019); Lu et al. (2019); Korba et al. (2020); Wang and Li (2020, 2022); Lu et al. (2023); Kook et al. (2024); Liu et al. (2024); Yan et al. (2024); Yao et al. (2024) and the references therein for further developments.
@@
In Section 2, we state the problem setup, and discuss some relevant properties of MFVI. In Section 3, we lay out the McKean–Vlasov equation corresponding to the MFVI problem, and introduce a particle algorithm. In Section 4, we state our main theorem which provides a convergence guarantee for the stochastic particle algorithm, Algorithm 2. In Section 5, we prove the theoretical results stated in Section 4. Some technical lemmas used in the proof are given in Section 6. Section 7 is devoted to discussions.
@@
We set up some notations that will be used throughout the paper. For any $d\in\mathbb{Z}_{+}=\{1,2,\cdots\}$, we denote $[d]=\{1,2,\cdots,d\}$ and let $\bm{I}_{d}$ represent the identity matrix of size $d$. For any $d\in\mathbb{Z}_{+}$, we denote by $\mathcal{P}(\mathbb{R}^{d})$ the set of probability distributions on $\mathbb{R}^{d}$, and by $\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d})$ the family of probability distributions on $\mathbb{R}^{d}$ that are absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{d}$. We denote the Euclidean norm by $\|\cdot\|_{2}$. For any two symmetric matrices $\bm{A}$ and $\bm{B}$ of the same size, we write $\bm{A}\preceq\bm{B}$ or $\bm{B}\succeq\bm{A}$ if $\bm{B}-\bm{A}$ is positive semidefinite. For any random variable $X$, we denote the law of $X$ by $\mathrm{Law}(X)$. For any $\bm{\mu}\in\mathbb{R}^{d}$ (where $d\in\mathbb{Z}_{+}$) and $d\times d$ positive definite matrix $\bm{\Sigma}$, we denote by $N(\bm{\mu},\bm{\Sigma})$ the multivariate normal distribution with mean vector $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$. For any two quantities $A,B>0$, we write $A=\mathcal{O}_{K}(B)$ if $A\leq CB$ for a positive constant $C$ that only depends on $K$. For any function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ (where $d\in\mathbb{Z}_{+}$), we denote its gradient and Hessian matrix by $\nabla f$ and $\nabla^{2}f$, respectively, and denote $\partial_{i}f(\bm{x}):=\frac{\partial f(\bm{x})}{\partial x_{i}}$ for any $\bm{x}=(x_{1},\cdots,x_{d})\in\mathbb{R}^{d}$ and $i\in[d]$.
@@ 2 Mean-field variational inference
In this section, we introduce the optimization problem involved in mean-field variational inference, and discuss relevant properties of this optimization problem.
@@ 2.1 Problem setup
Here $D_{\mathrm{KL}}(q\|p)=-\mathbb{E}_{q}\log(\frac{\mathrm{d}p}{\mathrm{d}q})$ is the Kullback–Leibler divergence (also known as the relative entropy) from $p$ to $q$. An optimal solution to (2.2) is called a mean-field approximation of $p$. MFVI is usually implemented using an algorithm called coordinate ascent variational inference (CAVI). In addition to the mean-field assumption, CAVI requires each complete conditional of the underlying model to belong to an exponential family (Bishop, 2006; Blei et al., 2017). To improve the expressive power of the variational family, nonparametric variants of CAVI have been proposed. For instance, $\mathcal{Q}$ can be the set of Gaussian mixtures (Gershman et al., 2012; Lambert et al., 2022). In this paper, we work with the entire mean-field family (2.1).

### v3

@@ Abstract
Variational inference is a fast and scalable alternative to Markov chain Monte Carlo and has been widely applied to posterior inference tasks in statistics and machine learning. A traditional approach for implementing mean-field variational inference (MFVI) is coordinate ascent variational inference (CAVI), which relies crucially on parametric assumptions on complete conditionals. We introduce a novel particle-based algorithm for MFVI, named PArticle VI (PAVI), for nonparametric mean-field approximation. We obtain non-asymptotic error bounds for our algorithm. To our knowledge, this is the first end-to-end guarantee for particle-based MFVI.
@@ 1 Introduction
Variational inference (VI) is a popular Bayesian inference method for approximating intractable posterior distributions (jordan1999introduction; wainwright2008graphical; blei2017variational). The approximation is obtained by solving a constrained optimization problem:
@@
where $\pi\in\mathcal{P}(\mathbb{R}^{n})$ is the posterior distribution, $\mathcal{Q}\subseteq\mathcal{P}(\mathbb{R}^{n})$ is some set of probability distributions, and $D_{\mathrm{KL}}(\cdot\|\cdot)$ is the Kullback–Leibler divergence. A common choice for $\mathcal{Q}$ is the set of product measures, i.e., $\mathcal{Q}=\big\{\prod_{i=1}^{n}q_{i},\;q_{i}\in\mathcal{P}(\mathbb{R})\big\}$, which is called the mean-field family. The corresponding variational problem for this choice of $\mathcal{Q}$ is called mean-field variational inference (MFVI). Compared to Markov chain Monte Carlo (MCMC) methods, which can have long burn-in periods, the optimization approach of VI allows for faster and more scalable inference in practice (blei2017variational).
@@
Optimization problems over spaces of probability distributions can in theory be solved with a gradient flow by endowing the space of probability distributions with the 2-Wasserstein metric $\mathcal{W}_{2}$ and using a discrete scheme (ambrosio2005gradient; santambrogio2017euclidean). The use of $\mathcal{W}_{2}$–gradient flows to solve the variational problem (1.1) has led to developments in theoretical understanding of VI (lambert2022variational; yao2022mean; diao2023forward; Lac23) as well as the unconstrained version of (1.1) whose solution is the exact posterior (garcia2020bayesian). Furthermore, a seminal work jordan1998variational introduces a close connection between $\mathcal{W}_{2}$–gradient flows and PDEs: the discrete iterative scheme that describes the $\mathcal{W}_{2}$–gradient flow converges, in the small step-size limit, to the solution to a Fokker–Planck equation (jordan1998variational, Theorem 5.1). The solution to the Fokker–Planck equation, in turn, describes the law of a particle evolving according to a stochastic differential equation (SDE). The differential equation characterization of the $\mathcal{W}_{2}$–gradient flow opens up avenues for computation and theoretical analysis.
@@
Since VI operates in a subset $\mathcal{Q}$ of the $\mathcal{W}_{2}$ space, efforts have been made to describe gradient flows for different choices of $\mathcal{Q}$. One challenge is to prove that the $\mathcal{W}_{2}$–gradient flow constrained to $\mathcal{Q}$ is still a gradient flow. When this is shown, theoretical guarantees follow (lambert2022variational; yao2022mean; Lac23). In the case of MFVI, a recent paper Lac23 constructs a $\mathcal{W}_{2}$–gradient flow constrained to the submanifold of product measures. This gradient flow is the solution to a nonlinear Fokker–Planck equation, which is a system of coupled PDEs. The SDE corresponding to the PDEs is a McKean–Vlasov equation obtained through independent projections of diffusions. The paper also specifies the convergence rate of the gradient flow to the MFVI solution.
@@
In this paper, we introduce a particle-based algorithm for MFVI, which we term PArticle VI (PAVI). The algorithm is built upon the SDE description of MFVI and stochastic approximation. Under standard regularity assumptions, we provide an upper bound on the 2-Wasserstein distance between the particle approximation and the solution to (1.1). Notably, our convergence guarantee is non-asymptotic and holds for any number of particles and total time duration. To our knowledge, this is the first end-to-end guarantee for particle-based MFVI.
@@
We outline recent theoretical advances in VI based on gradient flows in the $\mathcal{W}_{2}$ space as follows. When $\mathcal{Q}$ is the set of Gaussians, the resulting form of VI is called Gaussian VI. The set of Gaussians endowed with the $\mathcal{W}_{2}$ distance is also called the Bures–Wasserstein (BW) space (bures1969extension), which is a subset of the $\mathcal{W}_{2}$ space. By studying $\mathcal{W}_{2}$–gradient flows constrained to the BW space, lambert2022variational propose a stochastic gradient descent algorithm for Gaussian VI—which discretizes the gradient flow—and provide theoretical guarantees, and diao2023forward develop a forward-backward algorithm along with convergence rates for Gaussian VI. These works suggest extensions to the case when $\mathcal{Q}$ is the set of mixtures of Gaussians, which corresponds to an essentially nonparametric $\mathcal{Q}$. See yi2023bridging; huix2024theoretical for further developments.
@@
In the mean-field case where $\mathcal{Q}$ consists of product distributions, yao2022mean establish convergence guarantees for a coordinate-wise $\mathcal{W}_{2}$–gradient flow for MFVI in Bayesian models, and they propose two algorithms for practical implementation: a particle algorithm based on the Langevin SDE, and a function approximation method using neural networks. Their particle algorithm suffers from cumulative numerical errors that persist in the long-time limit, see Lemma H.1 and Theorem D.2 therein. Also, the error incurred by particle approximation is not analyzed. Lac23 shows that MFVI can be described by a $\mathcal{W}_{2}$–gradient flow constrained to the submanifold of product distributions and establishes continuous-time convergence guarantees. The difficulty in developing an efficient algorithm is also discussed. JCP23 propose the approximation of the MFVI solution through optimization over polyhedral subsets of the Wasserstein space, and introduces projected gradient descent algorithms with theoretical guarantees. Aside from those algorithmic and computational developments, there has also been recent work on the statistical accuracy of the solution to the MFVI problem for approximating the true posterior $\pi$. We refer the reader to (zhang2020convergence; mukherjee2022variational; lacker2024mean) and the references therein for details.
@@
Our convergence result resembles those from the literature on Langevin Monte Carlo (LMC), see dalalyan2017theoretical; durmus2017nonasymptotic; cheng2018convergence; durmus2019analysis; balasubramanian2022towards. LMC simulates an SDE whose stationary distribution is proportional to $e^{-V(\cdot)}$, where $V$ is a given potential. Non-asymptotic theoretical guarantees for LMC are given under the condition that the potential $V$ is continuously differentiable, strongly convex, and has a Lipschitz gradient. We refer the reader to chewi2024 for a survey.
@@
Our algorithm is built on the framework of gradient flows and particle algorithms. Gradient flows and particle algorithms have recently been applied to a wide range of problems in statistics and optimization. For example, saeedi2017variational propose a particle algorithm for approximating discrete distributions with the approximating family given by distributions with at most $K$ atoms (where $K$ is fixed). cheng2024particle propose particle algorithms for distribution approximation using generalized Wasserstein gradient flow of the Kullback–Leibler divergence. Their algorithms do not impose explicit specification of an approximating family $\mathcal{Q}$ and rely on neural networks to estimate certain vector fields used for updating. Their theoretical analysis relies on assumptions about neural network approximation accuracy and does not consider the error arising from particle approximation. We also refer the reader to liu2016stein; gallouet2017jko; chizat2018global; liu2019understanding; lu2019accelerating; korba2020non; wang2020information; wang2022accelerated; lu2023birth; kook2024sampling; liu2024towards; yan2024learning; YCY and the references therein for further developments.
@@
In Section 2, we state the problem setup and discuss some relevant properties of MFVI. In Section 3, we lay out the McKean–Vlasov equation corresponding to the MFVI problem, and introduce a particle algorithm. In Section 4, we state our main theorem, which provides a convergence guarantee for our algorithm. In Section 5, we prove the theoretical results stated in Section 4. Some technical lemmas used in the proof are given in Section 6. Section 7 is devoted to discussions.
@@
For any $d\in\mathbb{Z}_{+}=\{1,2,\cdots\}$, we denote $d]=\{1,2,\cdots,d\}$ and let $\bm{I}_{d}$ represent the identity matrix of size $d$. For any $d\in\mathbb{Z}_{+}$, we denote by $\mathcal{P}(\mathbb{R}^{d})$ the set of probability distributions on $\mathbb{R}^{d}$, and by $\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d})$ the family of probability distributions on $\mathbb{R}^{d}$ that are absolutely continuous with respect to the Lebesgue measure. We denote the Euclidean norm by $\|\cdot\|_{2}$. For any two symmetric matrices $\bm{A}$ and $\bm{B}$ of the same size, we write $\bm{A}\preceq\bm{B}$ or $\bm{B}\succeq\bm{A}$ if $\bm{B}-\bm{A}$ is positive semidefinite. For any random variable $X$, we denote its law by $\mathrm{Law}(X)$. For any $\bm{\mu}\in\mathbb{R}^{d}$ (where $d\in\mathbb{Z}_{+}$) and $d\times d$ positive definite matrix $\bm{\Sigma}$, we denote by $N(\bm{\mu},\bm{\Sigma})$ the multivariate normal distribution with mean vector $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$. For any two quantities $A,B>0$, we write $A=O_{K}(B)$ if $A\leq CB$ for a positive constant $C$ that only depends on $K$. For any function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ (where $d\in\mathbb{Z}_{+}$), we denote its gradient and Hessian matrix by $\nabla f$ and $\nabla^{2}f$, respectively, and denote $\partial_{i}f(\bm{x}):=\frac{\partial f(\bm{x})}{\partial x_{i}}$ for any $\bm{x}=(x_{1},\cdots,x_{d})\in\mathbb{R}^{d}$ and $i\in[d]$.
@@ 2 Mean-field variational inference
In this section, we introduce the optimization problem underlying mean-field variational inference and discuss its relevant properties.
@@ 2.1 Problem setup
Here $D_{\mathrm{KL}}(q\|p)=\mathbb{E}_{q}\log(\frac{\mathrm{d}q}{\mathrm{d}p})$ is the Kullback–Leibler divergence (also known as the relative entropy) from $p$ to $q$. As a convention, $D_{\mathrm{KL}}(q\|p)=+\infty$ if $q$ is not absolutely continuous with respect to $p$. An optimal solution to (2.2) is called a mean-field approximation of $p$. MFVI is usually implemented using coordinate ascent variational inference (CAVI) (bishop2006pattern; blei2017variational). In addition to the mean-field assumption, CAVI requires each complete conditional of the underlying model to belong to an exponential family. To improve the expressive power of the variational family, nonparametric variants of CAVI have been proposed. For instance, $\mathcal{Q}$ can be the set of Gaussian mixtures (gershman2012nonparametric; lambert2022variational). In this paper, we work with the entire mean-field family (2.1).
Remark 1 (Block MFVI).
Prior works on MFVI have also used products of multivariate distributions to form block mean-field families (bishop2006pattern). Given positive integers $\{d_{i}\}_{i=1}^{m}$ and a target distribution $p\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d_{1}+\cdots+d_{m}})$, block MFVI approximates $p$ by an optimal solution to the problem (2.2) with
$\mathcal{Q}=\bigg\{q^{1}\otimes\cdots\otimes q^{m}:~q^{i}\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d_{i}}),~~\forall i\in[m]\bigg\}\,.$ (2.3)
While our main discussion focuses on the mean-field family (2.1) for notational simplicity, we will also present algorithmic and theoretical results for the general scenario.